# Enhancing Biomedical Knowledge Discovery for Diseases: An Open-Source Framework Applied on Rett Syndrome and Alzheimer’s Disease

### v2

@@ 4 Results
Table 3: LaMReDA and LaMReDM Results (%) in binary and multi-class setup (BiomedBERT $\square$:base,$\blacksquare$:large): Each cell (unless cross-disease experiments) shows the average F1-score from 10 runs (original test set) and from 5-fold cross-validation setup.
@@
Human Performance. To assess and compare to human performance, two additional experts identify the relation type in a random sample of 300 instances from the test set of each dataset (Tab. 1). The evaluation ground truth is based on the original test set labels. In the binary setup, the average F1-score ranges from 92.14 for ReDReS to 91.87 for ReDAD. The LaMReDA, LaMReDM, and LaMEL models achieve performance comparable to human experts, indicating a high ability to detect semantic relations. Multi-class macro F1-scores range from 85.23 (micro: 85.45) to 85.76 (micro: 85.87) for ReDReS and ReDAD, respectively. Compared to human experts, all models show a performance gap, highlighting that identifying more complex aspects of semantic relation is a challenging task.
@@
Baseline performance - lower bound. We randomly assign labels based on the training data’s class distribution (Tab 1). In the binary setup, the baseline achieves F1-scores of 54% (ReDReS) and 53.16% (ReDAD). For the multi-class setup, the macro F1-scores range from 32.05% to 32.43%, stressing the task’s difficulty, particularly for distinguishing various semantic relations (multi-class) [^11] More information is available in Appendix F..
5 Probing
@@
This study probes BiomedBERT’s ability to capture semantic relations between entities. We explore different transformer layer representations and attention scores per layer and attention head. Averaged pooled entity representations are extracted from each layer, followed by training a linear classification layer. We test relation representations $R_{D}$, $R_{O}$, and $R_{P}$ (Eq. 12, 23, 24) of LaMReDA and LaMReDM to assess the impact of the context vector. Out-of-the-box representations are evaluated without the projection linear layer $l()$. We also extract average attention scores of tokens for each entity towards the other across each layer and head, concatenating these into a feature vector for training a linear classification layer. Following Chizhikova et al. (2022), we also train the classification layer using average attention scores between the two entities across all layers.
@@
Figure 4: ReDReS Probing (Binary setup): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Tab. 3). Second boundary: classifier with average attention scores across all layers as input.
@@
Fig. 4 shows the results of probing experiments in the binary setup using ReDReS and BiomedBERT base. [^12] Additional probing experiments in Appendix I. The 10th and 11th layers provide the most informative representations for relation types ($R_{D}$, $R_{O}$, and $R_{P}$). The $R_{D}$ representation, using element-wise multiplication, outperforms other representations in intra-layer comparisons, suggesting its effectiveness without end-to-end training. However, as highlighted in section 4, inter-model comparisons indicate that the transformer layers and projection layer $l()$ capture crucial information for relation detection, regardless of the aggregation function. Using context vectors with $R_{O}$ and $R_{P}$ generally offers no advantage, though $R_{O}$ from the 10th and 11th layers performs well, indicating possibly meaningful localized context. Attention scores between entities in the 12th layer yield the best performance, surpassing the baseline of using scores from all layers, indicating strong attention between the entities in the last layer. Figure 4 reveals that the 6th and 9th attention heads are most informative for relation detection.
6 Related Work
Information Extraction Datasets. Several biomedical datasets aim to enhance Information Extraction (IE) system development (Theodoropoulos et al., 2021; Nasar et al., 2021; Detroja et al., 2023; Huang et al., 2024), typically focusing on one or a few entity types and their interactions. AIMed (Bunescu et al., 2005), BioInfer (Pyysalo et al., 2007), and BioCreative II PPI IPS (Krallinger et al., 2008) formulate protein-protein interactions. The chemical-protein and chemical-disease interactions are modeled by DrugProt (Miranda et al., 2021) and BC5CDR (Li et al., 2016), respectively. ADE (Gurulingappa et al., 2012), DDI13 (Herrero-Zazo et al., 2013), and n2c2 2018 ADE (Henry et al., 2020) include drug-ADE (adverse drug effect) and drug-drug interactions. EMU (Doughty et al., 2011), GAD (Bravo et al., 2015) and RENET2 (Su et al., 2021) contain relations between genes and diseases. N-ary (Peng et al., 2017) incorporates drug-gene mutation interactions. The task of event extraction is illustrated by GE09 (Kim et al., 2009), GE11 (Kim et al., 2011), and CG (Pyysalo et al., 2013). DDAE (Lai et al., 2019) includes disease-disease associations. BioRED (Luo et al., 2022a) focuses on document-level relations for various entities. Unlike these datasets, ReDReS and ReDAD focus on RS and AD, include entities of up to 82 different semantic types, and model the semantic relation between them.
Knowledge Discovery. Gottlieb et al. (2011) present PREDICT, a method for ranking potential drug-disease associations to predict drug indications. Romano et al. (2024) release AlzKB, a heterogeneous graph knowledge base for AD, constructed using external data sources and describing various medical entities (e.g., chemicals, genes). Other graph-based efforts model knowledge around AD for tasks such as drug repurposing (Daluwatumulle et al., 2022; Nian et al., 2022; Hsieh et al., 2023), gene identification (Binder et al., 2022), or as general knowledge repositories (Sügis et al., 2019). Another paradigm for knowledge discovery is the open information extraction (OIE) setup (Etzioni et al., 2008; Mausam et al., 2012), which faces challenges such as data consistency, performance evaluation, and semantic drift (Zhou et al., 2022). Research efforts (Nebot and Berlanga, 2011; Movshovitz-Attias and Cohen, 2012; Nebot and Berlanga, 2014; de Silva et al., 2017; Wang et al., 2018) aim to address these problems and extract knowledge with little or no supervision. Advances in literature-based discovery (Gopalakrishnan et al., 2019; Thilakaratne et al., 2019) try to identify novel medical entity relations using graph-based (Kilicoglu et al., 2020; Nicholson and Greene, 2020), machine learning (Zhao et al., 2021; Lardos et al., 2022), and co-occurrence methods (Kuusisto et al., 2020; Millikin et al., 2023). Tian et al. (2024) stress the potential of large language models (LLMs) to summarize, simplify, and synthesize medical evidence (Peng et al., 2023; Tang et al., 2023; Shaib et al., 2023), suggesting that LLMs may have encoded biomedical knowledge (Singhal et al., 2023). To exploit this potential, we explore constructing LM representations for knowledge discovery. To the best of our knowledge, no systematic approach assembles knowledge about RS. Unlike previous work, we introduce a, in principle, disease-agnostic framework, to acquire knowledge about RS and AD starting from raw text.
7 Conclusion
@@
This work presents an open-source framework for disease knowledge discovery from raw text. We contribute two new annotated datasets for RS (ReDReS) and AD (ReDAD), facilitating further research. Extensive evaluation explores various methods for representing relations and entities, yielding insights into optimal modeling approaches for semantic relation detection, and emphasizing language models’ potential in knowledge discovery.
Limitations
One limitation of the paper is that the data pipeline relies on an external mention extractor/linker. However, this aspect introduces flexibility, allowing researchers and practitioners to integrate custom models suited to their specific applications. The creation of gold-standard datasets requires the manual work of medical experts. This process is time-consuming and resource-intensive, potentially limiting the scalability of the approach. Nevertheless, the experiments demonstrate that the supervised models of the study achieve strong performance in semantic relation detection without needing a large training set. Additionally, the cross-disease experiments highlight the robustness of the models in both binary and multi-class setups. This finding enables transfer learning scenarios in semantic relation detection, which can be applied to other diseases or medical aspects, indicating a potential for broader applications and research opportunities.
Ethics Statement
All recruited medical experts provided informed consent before participating in the annotation process. The compensation provided to the annotators was adequate and considered their demographic, particularly their country of residence.
@@ Appendix A Data Pipeline: Additional Information
To facilitate effective abstract retrieval, we implement an iterative approach to circumvent the API’s limitation of retrieving only 10,000 article IDs per query. This iterative process enables us to access a comprehensive set of PubMed IDs (PMIDs) related to the query. The detailed list of the 82 semantic types of the MetaMapLite-based pipeline is presented in Table 4. In addition to the MetaMapLite-based pipeline, we propose a second pipeline that is based on ScispaCy (Apache License 2.0) (Fig. 5). The difference lies in the selection of entity extractors and linkers that map the extracted entities to knowledge schemes. Unlike MetaMapLite, which adopts an integrated approach where mention extraction and linking are performed simultaneously in a single step and focuses on UMLS mapping, allowing for more precise and targeted extraction of entities, ScispaCy serves a broader range of Natural Language Processing (NLP) tasks. After the retrieval of the abstracts that is described in subsection 2.1, the following steps are executed: Knowledge schema and linker generation, Mention extraction, Entity linking and Sampling of linked identifiers.
@@
Knowledge schema and linker generation. ScispaCy (Neumann et al., 2019) harnesses an older version of UMLS (2020AA). This version serves as the foundation upon which ScispaCy trains and constructs its linkers that operate on a char-3grams string overlap-based search mechanism, facilitating efficient and accurate entity recognition and linking processes. Following the paradigm of ScispaCy, we provide scripts for generating updated linkers tailored to a range of knowledge schemes. These include UMLS (Bodenreider, 2004), Gene Ontology (GO) (Consortium, 2004), National Center for Biotechnology Information (NCBI) taxonomy (Schoch et al., 2020), RxNorm (Nelson et al., 2011), SNOMED Clinical Terms (SNOMEDCT_US) (Stearns et al., 2001), Human Phenotype Ontology (HPO) (Köhler et al., 2021), Medical Subject Headings (MeSH) (Lipscomb, 2000) DrugBank (Knox et al., 2024) and Gold Standard Drug Database (GS) [^13] https://www.nlm.nih.gov/research/umls/
sourcereleasedocs/current/GS/index.html
. Of particular note is the inclusion of UMLS, a unified system encompassing various knowledge bases, vocabularies, taxonomies, and ontologies pertinent to the biomedical domain. Any supported linker maps the concepts to UMLS CUIs enhancing the standardization of medical terminology. Notably, the flexibility of ScispaCy’s implementation allows for seamless expansion to incorporate additional knowledge bases, thereby enhancing its versatility and applicability across diverse research needs.
@@
Mention extraction. ScispaCy boasts four distinct entity extractors, each trained on different corpora, collectively encompassing a range of entity types. These extractors include named entity recognition (NER) models trained on the CRAFT corpus (with 6 entity types) (Bada et al., 2012), JNLPBA corpus (with 5 entity types) (Collier et al., 2004), BC5CDR corpus (with 2 entity types) (Li et al., 2016), and BIONLP13CG corpus (with 16 entity types) (Kim et al., 2013). To maximize the range of the entity extraction, we leverage these diverse extractors in tandem, allowing us to capture mentions of 18 unique entity types (gene or protein, cell, chemical, organism, disease, organ, DNA, RNA, tissue, cancer, cellular component, anatomical system, multi-tissue structure, organism subdivision, developing anatomical structure, pathological formation, organism substance, and immaterial anatomical entity).
@@
Entity linking. This process enhances the semantic understanding of the extracted entities, facilitates standardization, which is a key challenge in the biomedical field (Bettencourt-Silva et al., 2012; Theodoropoulos et al., 2023), and promotes interoperability with external resources by associating the entities with specific concepts in supported knowledge schemes. Each entity is subjected to a linking process where we attempt to map it to concepts within supported knowledge bases or vocabularies. If a match is found, the entity is assigned a unique identifier, referred to as a CUI, corresponding to the specific concept in the knowledge schema. As entities may be linked to multiple knowledge sources, we merge the extracted CUIs obtained from the different linkers. This consolidation process ensures that each entity is associated with a comprehensive set of identifiers, encompassing diverse perspectives and representations across various knowledge schemes.
@@
The co-occurrence graph generation step is described in subsection 2.1.
Semantic Types
Bacterium Body Substance Body System Behavior Biologic Function
Plant Organism Population Group Receptor Reptile
Virus Vitamin Vertebrate
Table 4: List of the 82 semantic types of the MetaMapLite-based pipeline.
@@
Algorithm 1 Sentence Sampling.
$sent\_c$, $co\_g$, $n$
Initialize dictionary $f\_d$
for $s\_id$, $sent$ in $sent\_c$ do
     $c\_p\leftarrow extract\_conc(sent)$
     $f\_p\leftarrow extract\_freq(c\_p,co\_g)$
     $t\_f\leftarrow sum(f\_p)$
     $f\_d\leftarrow save(s\_id,c\_p,f\_p,t\_f)$
end for
Initialize lists $ids$, $f\_l$, $inv\_f\_l$
for $s\_id$ in $f\_d$ do
     $ids\leftarrow append(s\_id)$
     $t\_f\leftarrow get(f\_d,s\_id)$
     $f\_l\leftarrow append(t\_f)$
     $inv\_f\_l\leftarrow append(1/t\_f)$
end for
$t\_f\_sum\leftarrow sum(f\_l)$
$inv\_t\_f\_sum\leftarrow sum(inv\_f\_l)$
Initialize lists $prob$, $inv\_prob$
for $f$ in $f\_l$ do
     $p=f/t\_f\_sum$
     $prob\leftarrow append(p)$
end for
for $f$ in $inv\_f\_l$ do
     $p=f/inv\_t\_f\_sum$
     $inv\_prob\leftarrow append(p)$
end for
$\mathcal{P}\leftarrow prob\_distr(ids,prob)$
$\mathcal{IP}\leftarrow prob\_distr(ids,inv\_prob)$
$sam\_sent\_1\leftarrow sample(\mathcal{P},n/2)$
$sam\_sent\_2\leftarrow sample(\mathcal{IP},n/2)$ return $sam\_sent\_1$, $sam\_sent\_2$

### v3

@@ IV-D Experiments with Additional Language Models
    Base version.
V Results
Tables IV and VI report the F1-scores for the LaMReDA, LaMReDM, and LaMEL models on the ReDReS and ReDAD datasets. Each cell (except for the cross-disease experiments) displays two values: the average F1-score from 10 runs on the original test set (Table II) and the average F1-score from a 5-fold cross-validation. The models perform well across all relation (A-P) and entity (A-H) representations, showing their ability to learn meaningful representations for the semantic relation task regardless of initial token selection. However, we observe patterns regarding relation representations. In the binary setup, the relation representation $R_{G}$ (15) yields strong results for both datasets, suggesting that including the [CLS] token representation might be beneficial. In the multi-class setup, relation representations $R_{L}$ (20), $R_{J}$ (18), and $R_{O}$ (23) are effective for both datasets, indicating that the surrounding context is crucial for the more complex task, as $R_{L}$ and $R_{J}$ include the averaged pooled representation of intermediate tokens between entities, and $R_{O}$ leverages the context vector [81]. The intra-model comparison reveals that over-parameterization tends to be useful. Using BiomedBERT large generally results in better performance than the base alternative. The BiomedBERT base shows superior performance mainly only in experiments using the original splits of ReDReS (Table II). LaMEL is highly competitive with LaMReDA and LaMReDM, indicating that learning entity embedding spaces optimized for relation detection is promising. LaMEL achieves the highest performance in the 5-fold setup of ReDAD and the original setup of ReDReS, with F1-scores of 91.03% and 91.25% respectively.
The inter-model comparison across the same relation representations indicates that the aggregation function does not significantly impact the relation detection task. Neither LaMReDA (element-wise addition) nor LaMReDM (element-wise multiplication) show a clear advantage over the other. This suggests that the transformer layers of BiomedBERT and the projection layer $l()$ preceding the aggregation are effectively trained in both models to encode the essential information for relation detection, regardless of the aggregation function used. The cross-disease experiments underscore the robustness of the models in both binary and multi-class setups. This robustness supports transfer learning [87] in semantic relation detection, extending to other diseases, highlighting the potential for broader applications and research endeavors in knowledge discovery.
@@ V Results
Human Performance. To assess and compare the human performance, two additional experts identify the relation type in a random sample of 300 instances from the test set of each dataset (Table II). The evaluation ground truth is based on the original test set labels. In the binary setup, the average F1-score ranges from 92.14 for ReDReS to 91.87 for ReDAD. The LaMReDA, LaMReDM, and LaMEL models achieve performance comparable to those of human experts, indicating a high ability to detect semantic relations. Multi-class macro F1-scores range from 85.23 (micro: 85.45) to 85.76 (micro: 85.87) for ReDReS and ReDAD, respectively. Compared with human experts, all models show a performance gap, highlighting that identifying more complex aspects of semantic relations is a challenging task.
@@
Baseline performance - lower bound. Labels are randomly assigned labels based on the class distribution of the training data (Table II). In the binary setup, the baseline achieves F1-scores of 54% (ReDReS) and 53.16% (ReDAD). For the multi-class setup, the macro F1-scores range from 32.05% to 32.43%, stressing the task’s difficulty, particularly for distinguishing various semantic relations (multi-class). Appendix E provides additional information.
The results presented in Tables V and VII demonstrate the F1-scores of the LaMReDA and LaMEL models on the ReDReS and ReDAD datasets, using different language models. The experiments show that both LaMReDA and LaMEL perform consistently well across various relations ($R_{A}$, $R_{D}$, $R_{E}$, $R_{F}$, $R_{G}$, $R_{O}$) and entities ($E_{A}$, $E_{D}$, $E_{E}$, $E_{H}$) representations, utilizing the different language models. This indicates the models’ ability to learn meaningful representations for the semantic relation detection task, independent of the initial token selection.
An intra-lm comparison reveals that the use of relation representation $R_{O}$ results in lower performance across both datasets when BioGPT is employed, particularly in the multi-class setup (Table V). This suggests that the context vector extracted from the generative model is less informative for this task. Conversely, when the context vector is derived from BiomedBERT and BioLinkBERT, the models show strong performance, showing the best results on ReDAD in the 5-fold cross-validation setup (Table V).
The inter-language model comparison indicates that BiomedBERT and BioLinkBERT perform similarly, suggesting that leveraging document links during pre-training, as BioLinkBERT does, does not offer a significant advantage for our task. In contrast, the use of BioGPT leads to lower performance, particularly in the multi-class setup, emphasizing the superiority of encoder-based language models such as BiomedBERT and BioLinkBERT over decoder-based BioGPT for the relation detection task. Despite BioGPT’s larger size, with 347 million parameters [84] compared with the 110 million parameters of BiomedBERT and BioLinkBERT [10, 83], the generative approach appears less effective for this particular semantic relation task, further underscoring the advantages of encoder-focused models in this context.
Figure 6: ReDReS Confusion Matrices for each relation class using the best-performing model on the ReDReS dataset based on the macro evaluation (Table IV), with BiomedBERT-base as the backbone language model.
Figure 7: ReDReS Confusion Matrices for each relation class using the best-performing model on the ReDReS dataset based on the macro evaluation (Table IV), with BiomedBERT-large as the backbone language model.
V-A Error Analysis
Whereas the macro and micro evaluations in the multi-class setup indicate strong overall performance across different relation types, we conduct an error analysis to provide additional qualitative insights. Using the best-performing models on the ReDReS and ReDAD datasets based on the macro evaluation (Table IV), we extract predictions from the 5-fold cross-validation setup. Confusion matrices (tables that are used to analyze the classification performance, by displaying the number of true positives, true negatives, false positives, and false negatives) for each class are generated to present a comprehensive view of the misclassified instances (Fig. 6, 7, 8, 9). For example, the confusion matrix for the positive class is presented in the top-left side of Fig. 6, where the number of true positives, true negatives, false positives, and false negatives are 1344, 3255, 274 and 386, respectively.
Figure 8: ReDAD Confusion Matrices for each relation class using the best-performing model on the ReDAD dataset based on the macro evaluation (Table IV), with BiomedBERT-base as the backbone language model.
Figure 9: ReDAD Confusion Matrices for each relation class using the best-performing model on the ReDAD dataset based on the macro evaluation (Table IV), with BiomedBERT-large as the backbone language model.
We analyze the false negatives for each relation class to gain insights into the misclassification rates. In the ReDReS dataset, the majority of misclassified instances of the positive relation (324 and 331 instances using BiomedBERT-base and BiomedBERT-large, respectively) are labeled as complex relations (Table VIII). This is intuitively reasonable, as the model recognizes that the entities are semantically related but fails to predict the correct relation type. Conversely, complex relation instances are often mislabeled as no relation (Table VIII), highlighting the challenge of detecting nuanced semantic relations that require deeper reasoning. For the no relation class (Table VIII), most false negatives (364 and 303 instances using BiomedBERT-base and BiomedBERT-large, respectively) are also labeled as complex relations, indicating a model bias towards predicting associations between co-occurring entities. Similar patterns are observed in the ReDAD dataset regarding the positive, complex, and no relation classes. The under-represented negative class has the most misclassified cases labeled as complex or positive relations in ReDReS and ReDAD (Table VIII), respectively, suggesting that while the model detects the semantic connection, it tends to predict one of the two more represented classes in the datasets.
Different strategies can be applied during the training phase to improve model performance to address the class imbalance in the datasets. In the multi-class setup, the classes such as positive, complex, and no relation are over-represented. We can handle this imbalance by undersampling these over-represented classes, reducing their instances to approximate the frequency of the under-represented negative class. This approach ensures that the model does not become biased toward the majority classes, encouraging it to learn from a more balanced distribution of samples. However, due to the limited number of under-representative class instances, oversampling of negative class by incorporating additional samples from the distantly supervised datasets may be a more effective strategy. By increasing the number of instances of the negative class, the model may learn its distinct characteristics, thereby enhancing the classification performance. Another approach to managing class imbalance is to adjust the loss function by increasing the weight of the under-represented class. By assigning a higher rescaling weight to this class during loss calculation, the model is penalized more heavily for misclassifying instances from the negative class. This encourages the model to focus on improving its performance on this less frequent class, balancing the impact of each class on the overall learning process.
TABLE VIII: False negative analysis for the Positive, Complex, Negative, and No Relation relation class.
Data Class BiomedBERT-base BiomedBERT-large
Positive Complex Negative No Relation Positive Complex Negative No Relation
ReDReS Positive - 331 14 41 - 324 11 42
Complex 221 - 26 307 188 - 8 249
Negative 25 28 - 1 15 34 - 1
No Relation 28 303 2 - 41 364 0 -
ReDAD Positive - 424 27 40 - 425 21 18
Complex 335 - 19 338 274 - 6 297
Negative 44 23 - 2 44 25 - 1
No Relation 52 653 1 - 50 607 1 -
VI Distantly Supervised Datasets
ReDReS and ReDAD include gold annotations for a small fraction of the extracted sentences. The pre-processed text consists of 28,622 and 1,301,429 additional sentences related to RS and AD respectively, without annotations of the semantic relation between the detected entities. Observing that the supervised models achieve performance levels comparable to those of human experts, we leverage the best-performing models to generate silver labels for the unannotated instances. For the binary setup, we employ:
  • LaMReDA (BiomedBERT large) with the relation representation $R_{A}$ (9) for the RS corpus.
  • LaMReDA (BiomedBERT large) with the relation representation $R_{J}$ (18) for the AD corpus.
For the multi-class setup, we use:
  • LaMReDA (BiomedBERT large) with the relation representation $R_{F}$ (14) for the RS corpus.
  • LaMReDA (BiomedBERT large) with the relation representation $R_{F}$ (14) for the AD corpus.
We stress that the selection of the models relies on the performance in the 5-fold cross-validation setup to avoid choosing based on the model performance on the original test set. Each model is trained 10 times with different seeds using the original splits (Table II) of ReDReS and ReDAD. The best model weights are saved based on the performance on the development set. Every trained model provides predictions for the unannotated instances and the final silver labels are extracted through majority voting. The Distantly Supervised Relation Detection dataset for Rett Syndrome (DiSReDReS) contains 304,008 instances with 8,611 unique CUIs and 80 semantic types (Tables IX and X). The Distantly Supervised Relation Detection dataset for Alzheimer’s Disease (DiSReDAD) comprises 13,608,175 instances with 53,750 unique CUIs and 82 semantic types (Tables IX and X). As noisy labeling is inevitable in distantly supervised data and imposes challenges for knowledge extraction scenarios, the two extensive datasets can promote weakly supervised learning.
Weakly Supervised Setup. The task formulation remains the same as that described in Section IV. The train sets of ReDReS and ReDAD are replaced by DiSReDReS and DiSReDAD, respectively. The development and test sets are identical (Table II). To provide a benchmark, we train the LaMReDA (BiomedBERT base) with the relation representation $R_{A}$ (9) for 10 epochs utilizing the ADAM optimizer with learning rate 10-5. The batch is set to 32. The experiments are repeated 10 times with different seeds (42, 3, 7, 21, 77, 24, 69, 96, 44, 11) and the best scores are retained based on the performance on the development set.
TABLE IX: DiSReDReS and DiSReDAD: Statistics and performance (%)
Data Sentences Instances CUIs1 S.T.1 Benchmark3
Binary Multi-Class
Micro Macro
DiSReDReS 28,622 304,008 8,611 80 91.22±0.3 74.54±0.8 74.65±0.9
DiSReDAD 1,301,429 13,608,175 53,750 82 89.40±0.3 80.79±0.3 80.81±0.4
  • 1
    The total number of unique CUIs.
  • 2
    Semantic Types.
  • 3
TABLE X: DiSReDReS and DiSReDAD: Label Distribution
Dataset Labels - Type of relation
Positive Complex Negative No Relation
DiSReDAD 4,468,110 (32.8%) 5,755,884 (42.3%) 120,267 (0.9%) 3,263,914 (24.0%)
In the supervised setup (Table II), LaMReDA (BiomedBERT base) with the $R_{A}$ representation achieves 90.72% and 88.31% F1-score in the binary setup on ReDReS and ReDAD, respectively (Table IV). The multi-class macro F1-scores range from 74.52% (micro: 74.49%) to 77.34% (micro: 77.64%) for ReDReS and ReDAD, accordingly (Table IV). Table IX presents the benchmark performance in the binary and multi-class setup for both datasets. Notably, the performance is improved in the weakly supervised setup, indicating the robustness of LaMReDA, when trained with noisy data, and highlighting the quality of the silver labels of DiSReDReS and DiSReDAD.
VII Probing
@@
This study probes BiomedBERT’s ability to capture semantic relations between entities. We explore different transformer layer representations and attention scores per layer and attention head. The averaged pooled entity representations are extracted from each layer, followed by training a linear classification layer. We test relation representations $R_{D}$ (12), $R_{O}$ (23), and $R_{P}$ (24) of LaMReDA and LaMReDM to assess the impact of the context vector. Out-of-the-box representations are evaluated without the projection linear layer $l()$. We also extract the average attention scores of tokens for each entity towards the other across each layer and head, concatenating them into a feature vector for training a linear classification layer. Following [88], we also train the classification layer using the average attention scores between the two entities across all layers.
@@
Figure 10: ReDReS Probing (Binary setup) (BiomedBERT base): Examines LaMReDA/LaMReDM relation representations ($R_{D}$, $R_{O}$, $R_{P}$) and attention scores from each layer and explores average attention scores of tokens corresponding to each entity towards the other entity across attention heads. Top boundary: best LaMReDA and LaMReDM performance (Table IV). Second boundary: classifier with average attention scores across all layers as input.
@@
Fig. 10 shows the results of probing experiments in the binary setup using ReDReS and BiomedBERT base. The 10th and 11th layers provide the most informative representations for relation types ($R_{D}$, $R_{O}$, and $R_{P}$). The $R_{D}$ representation, using element-wise multiplication, outperforms other representations in intra-layer comparisons, suggesting its effectiveness without end-to-end training. However, as highlighted in Section V, inter-model comparisons indicate that the transformer layers and projection layer $l()$ capture crucial information for relation detection, regardless of the aggregation function. Using context vectors with $R_{O}$ and $R_{P}$ generally offers no advantage, though $R_{O}$ from the 10th and 11th layers performs well, indicating possibly meaningful localized context. Attention scores between entities in the 12th layer yield the best performance, surpassing the baseline of using scores from all layers, indicating strong attention between the entities in the last layer. Fig. 10 reveals that the 6th and 9th attention heads are most informative for relation detection. Additional probing experiments are presented in Appendix F.
VIII Conclusion
@@
This paper presents an open-source framework for disease knowledge discovery from raw text. Facilitating further research, we contribute two new annotated datasets [^6] All the recruited medical experts provided informed consent before participating in the annotation process. The compensation provided to the annotators was adequate, considering their demographic, particularly their country of residence. for RS (ReDReS) and AD (ReDAD) and the respective distantly supervised datasets DiSReDReS and DiSReDAD to promote weakly supervised scenarios. Extensive evaluation explores various methods for representing relations and entities, yielding insights into optimal modeling approaches for semantic relation detection, and emphasizing language models’ potential in knowledge discovery. In future work, additional existing and future LLMs can be evaluated for the task proposed in this paper.
@@ Appendix A Data Pipeline: Additional Information
In addition to the MetaMapLite-based pipeline, we propose a second pipeline based on ScispaCy (Apache License 2.0) (Fig. 11). The difference lies in the selection of entity extractors and linkers that map the extracted entities to the knowledge schemes. Other named entity recognition (NER) and linking tools, such as MedCAT [89], can also be incorporated. Unlike MetaMapLite, which adopts an integrated approach in which mention extraction and linking are performed simultaneously in a single step and focuses on UMLS mapping, allowing for more precise and targeted extraction of entities, ScispaCy serves a broader range of Natural Language Processing (NLP) tasks. After the retrieval of the abstracts described in Section III, the following steps are executed: Knowledge schema and linker generation, Mention extraction, Entity linking and Sampling of linked identifiers.
@@
Knowledge schema and linker generation. ScispaCy [90] harnesses an older version of UMLS (2020AA). This version serves as the foundation upon which ScispaCy trains and constructs its linkers that operate on a char-3grams string overlap-based search mechanism, facilitating efficient and accurate entity recognition and linking processes. Following the paradigm of ScispaCy, we provide scripts for generating updated linkers tailored to a range of knowledge schemes. These include UMLS [65], Gene Ontology (GO) [91], National Center for Biotechnology Information (NCBI) taxonomy [92], RxNorm [93], SNOMED Clinical Terms (SNOMEDCT_US) [94], Human Phenotype Ontology (HPO) [95], Medical Subject Headings (MeSH) [96] DrugBank [97] and Gold Standard Drug Database (GS) [^7] https://www.nlm.nih.gov/research/umls/sourcereleasedocs/current/GS/index.html. Of particular note is the inclusion of UMLS, a unified system encompassing various knowledge bases, vocabularies, taxonomies, and ontologies pertinent to the biomedical domain. Any supported linker maps the concepts to UMLS CUIs enhancing the standardization of medical terminology. Notably, the flexibility of ScispaCy’s implementation allows seamless expansion to incorporate additional knowledge bases, thereby enhancing its versatility and applicability across diverse research needs.
@@
Mention extraction. ScispaCy boasts four distinct entity extractors, each trained on different corpora, collectively encompassing a range of entity types. These extractors include NER models trained on the CRAFT corpus (with 6 entity types) [98], JNLPBA corpus (with 5 entity types) [99], BC5CDR corpus (with 2 entity types) [22], and BIONLP13CG corpus (with 16 entity types) [100]. To maximize the range of the entity extraction, we leverage these diverse extractors in tandem, allowing us to capture mentions of 18 unique entity types (gene or protein, cell, chemical, organism, disease, organ, DNA, RNA, tissue, cancer, cellular component, anatomical system, multi-tissue structure, organism subdivision, developing anatomical structure, pathological formation, organism substance, and immaterial anatomical entity).
@@
Entity linking. This process enhances the semantic understanding of the extracted entities, facilitates standardization, which is a key challenge in the biomedical field [101, 102], and promotes interoperability with external resources by associating the entities with specific concepts in the supported knowledge schemes. Each entity is subjected to a linking process where we attempt to map it to concepts within the supported knowledge bases or vocabularies. If a match is found, the entity is assigned a unique identifier, referred to as a CUI, corresponding to the specific concept in the knowledge schema. As entities may be linked to multiple knowledge sources, we merge the extracted CUIs obtained from the different linkers. This consolidation process ensures that each entity is associated with a comprehensive set of identifiers, encompassing diverse perspectives and representations across various knowledge schemes.
@@
The co-occurrence graph generation step is described in Section III.